# Securing Automated Insulin Delivery Systems: A Review of Security Threats and Protectives Strategies

### v1

@@ 4 Attack Vectors in AID systems
Redundant systems
Ransomware Attacks [60] The malware encrypts the system’s data or locks users out, demanding a ransom to restore access Jamming Attacks [57] Disrupting the wireless communication channels that the APS uses to operate Firmware/ Software Corruption [60] Corrupting the firmware or software that the APS relies on to function Routing protocol information attacks [61] Attacks on routing protocols to disrupt or manipulate the flow of information
Table 6: Attack vectors and common countermeasures in AID systems.
4.1 Threats to Confidentiality
Ensuring the confidentiality of diabetes devices aims to guarantee that sensitive information is accessible only to authorized users or systems. Threats to this security aspect can originate from unauthorized individuals or even from patients themselves.
4.1.1 Eavesdropping
This refers to the unauthorized interception of wireless communication which can expose (1) personal patient information related to therapy and physical condition, (2) device information, and (3) the device’s personal identification number (PIN). Laboratory experiments have shown that data transmitted by a commercial insulin delivery system can be intercepted using off-the-shelf Universal Software Radio Peripheral (USRP) and free software [12]. By capturing its identification number, attackers could then gain full control of the pump. At a 2012 security conference, an engineer showed a fatal dose of insulin could be delivered from 300 feet away [62]. These attacks exploit wireless communication vulnerabilities, such as wide transmission range and inadequate authentication and encryption. To mitigate these risks, countermeasures proposed in the literature mainly focus on (1) enhancing data protection through secure authentication, cryptography, and communication protocol, and (2) limiting transmission range by employing proximity communication channels.
4.1.2 Do-it-yourself (DIY) hacking
This refers to unauthorized access or modification of commercial diabetes devices by patients themselves driven by the desire for better data visualization and more advanced control with shorter official censor and approval time [16]. For instance, Nightscout is an open-source project that helps patients hack their Dexcom CGMs for remote glucose level monitoring on mobile devices. Projects such as DIYPS and #OpenAID systems enable patients to control insulin delivery. While these initiatives empower patients, they introduce risks due to vulnerabilities in unauthorized software and users’ potential lack of expertise. Compared to FDA-regulated medical software, these systems lack rigorous testing and responsible parties to support them in case of problems. To balance the need for patients to have greater control with secure AID systems against malicious exploitation, it is crucial to implement robust authentication strategies.
4.2 Threats to Integrity
The integrity of AID systems ensures that its data, algorithms, and operations remain accurate and consistent. Integrity can be compromised through two key aspects: data manipulation and algorithm manipulation.
4.2.1 Manipulation of Data
Attackers can manipulate wirelessly transmitted data, such as CGM readings or controller commands, to mislead the control algorithm or insulin pumps to issue incorrect insulin delivery. Data manipulation can occur through the following attack vectors:
  • 1.
    Replay Attacks: These attacks involve intercepting and replaying legitimate data packets to manipulate insulin delivery [6]. This type of attack is relatively simple to launch as it does not require device access and message decryption. Once the attacker gains access to the device (e.g., access the PIN), they can craft forged data packets to be accepted by the device [12], enabling bias injection and false data injection attacks.
  • 1.
    Bias Injection Attacks: In this attack, the attacker skews sensor readings by injecting a bias [39]. By making subtle, consistent changes, the attacker can cause the system to under or over-deliver insulin without needing extensive knowledge of the patient model or system operations. They can be difficult to detect due to their subtlety and correlation with the original data [63].
  • 1.
    False Data Injection Attacks (FDIA): FDIA could be categorized into random and targeted types. Random FDIA introduces arbitrary interference in the input data. Experiments in [56] showed that such attacks can cause outputs of ML-based glucose prediction models to change 10-20 mg/dL. This type is easier to detect through sequential change detection systems. In contrast, targeted FDIA synthesize precise input values to evade alarm and cause targeted harm to patients. Launching such attacks requires knowledge of the model’s architecture and parameters, but this information is often detailed in documentation. Kashyap et al. [56] formulated this attack as an optimization problem, demonstrating that an effective targeted FDIA with minimal, hard-to-detect perturbations can be computed within 60 milliseconds.
Defence strategies for these data manipulation attacks typically include developing intrusion detection systems to detect abnormal patterns in sensor data. Another defensive measure involves using redundant sensors or software monitors to cross-check data sources for inconsistencies.
4.2.2 Manipulation of Algorithms
Algorithm manipulation involves tampering with the computational logic or parameters setting of AID systems, targeting either control algorithms or physiologic models used for patient state prediction [41]. These include both traditional rule-based algorithms and machine learning-based models.
  • 1.
    Computational code attacks: These attacks compromise the command computational code in rule-based algorithms [64]. For example, Aliabadi et al. [65] identified that the synchronization between CGM and controller could be manipulated by altering the timing function (get-data-timer()). This modification allows attackers to delay or drop critical messages. Additional threats include altering essential settings, such as disabling pump alarms, modifying the maximum dosage limits, or enabling insulin overdoses [40][66][59].
  • 2.
    ML model attacks: ML models are increasingly used in AID systems to enhance performance [67]. For instance, the FDA-approved One Drop Blood Glucose Monitoring System employs a multi-layer perceptron (MLP) model for diabetes management. Deep reinforcement learning (RL) approaches have been explored for use in AID systems for BG control [68]. However, due to their black-box and data-dependent nature, understanding and monitoring the reasoning behind these models is challenging. Attackers may influence the model’s behaviour during training or inference phases or reverse-engineer the model to extract confidential information [54]. Moreover, ML-based algorithms can be severely interfered by a very slight input data perturbation that is hard to detect. In [56], minimal interference was calculated to misguide the ML model into predicting a patient’s glucose level as hyperglycaemic, leading the controller to administer an excessive dose of insulin. Experiments in [69] show that a small perturbation around ±1mg/dL on BG reading can lead to severe patient harm over time in an RL controller model.
To defend against algorithm integrity threats, formal verification techniques are often employed to assess the operations of rule-based algorithms. For ML-enabled systems, which pose greater challenges for formal verification, patient-specific insulin injection patterns have been modelled to detect abnormalities and inconsistencies.
4.3 Threats to Availability
  • 1.
    Denial-of-Service (DoS) and Distributed DoS (DDoS): In these attacks, an adversary floods the system with excessive traffic, overwhelming AID systems and preventing legitimate requests from being processed promptly. This could lead to delays or complete failure in delivering insulin.
  • 2.
    Jamming Attacks: By interfering with the wireless communication frequency, attackers can interrupt data flow, causing the system to fail to deliver insulin as required.
  • 3.
    Firmware/Software Corruption: Malicious firmware or software corruption can cause AID systems devices to malfunction or stop functioning entirely. Attacks exploiting software vulnerabilities or delivering harmful updates can prevent systems from operating as intended, leading to a cessation of insulin administration.
  • 4.
    Routing Protocol Information Attacks: Manipulating data transmission pathways in networked AID systems, attackers can disrupt or delay communication between key system components, compromising timely insulin delivery and glucose monitoring.
  • 5.
    Ransomware Attacks: Ransomware can encrypt data and lock critical components of AID systems, demanding ransom payments to restore access. This can prevent patients or healthcare providers from viewing essential data, such as glucose levels or delivery history, rendering the system inoperable until resolved.
Mitigation strategies for these threats include network traffic monitoring, secure communication protocols, authentication mechanisms, and the application of formal verification methods.
5 Defence Strategies for AID systems
Addressing the wide range of potential cyberattacks on AID systems requires a comprehensive defence strategy that spans multiple system layers. Defence mechanisms in AID systems can be grouped into three primary areas: protected communication channels, intrusion detection, and control activities monitoring. Besides these internal defences, external strategies—such as the use of redundant hardware or software components—are employed to detect abnormal activities independent of AID devices. Table 7 provides an overview of the defence types and corresponding methods proposed in past research for protecting AID systems.
Defence type Description Methods Protected communi-cation Authentication Verify the identity of users or devices Fingerprint recognition[70][71];
Voiceprint analysis[72];
Physical Unclonable Functions[73][74];
Blockchain[73][75]
Cryptography Transform information into an unreadable format for unauthorized users Rolling-code encryption[12];
AES-based encryption[76][57];
Asymmetric encryption [70];
Homomorphic encryption [77]
Secure communication protocols Define rules for secure data transmission RF-Harvested energy-based protocol[60];
Emergencies-resilient protocol[37];
Improved the Dexcom G4 protocol[57];
Mutual authentication protocol[78]
Proximity communication channels Require close physical proximity between the communicating devices Body-coupled communication[12];
Visible light communication[79]
Intrusion detection systems Signature-based Compare the real-time behaviour of the system against known security attacks Define computational, data integrity, and communication attacks[41];
Define abnormal signal’s physical and content features[5];
Define replay and DoS attacks[55];
Infusion site failures [80]; Pressure-induced sensor attenuations [81]
Specification-based Define specifications of AID systems’ normal operations and deem any violation as an intrusion Formal specification [82] with
personalized thresholds[83][59];
statistical analysis[84];
declarative property [41];
outlier detection[34][85];
specification mining[86];
external and internal state[87][88];
insulin predictive model[59]
Anomaly-based Identify data patterns that have different data characteristics from normal instances Model-based methods:
@@ 5 Defence Strategies for AID systems
Control strategies assessment Compare the dosage with the identified patient’s insulin infusion pattern Rule-based [41];
Data-driven[66][11];
Hybrid[83]
Redundant hardware or software components Use of multiple, often independent, components that perform the same function Hardware:
ECG monitor[88]; Universal Software Radio Peripheral[5]; Separate computation core[60]
Software:
@@
Replay attack detection filters[6]; Auxiliary safety layer for controllers[111]
Table 7: Defence strategies proposed for AID systems.
5.1 Protected Communication
@@ 5.1.1 Authentication Methods
Biometric authentication methods like fingerprint recognition [70][71] and voiceprint analysis [72] are user-friendly and offer strong uniqueness but introduce concerns about sensitive biometric data storage, which could be spoofed through advanced attacks. To mitigate these risks, hardware-based security methods such as Physically Unclonable Functions (PUFs) exploit inherent and unique physical variations in manufacturing processes, offering low computational overhead and an unclonable security mechanism without the need for key storage [73][74]. Additionally, blockchain technology has been explored to protect patient data by decentralizing data storage in central cloud servers [73]. Rahmadika et al. [75] further incorporated blockchain with federated learning for sharing local parameters for the misbehaviour detection model in AID systems without compromising patient privacy.
5.1.2 Encryption Techniques
Following successful authentication, establishing encrypted communication between AID systems components is vital for maintaining data confidentiality and integrity. A key challenge is the extreme energy and computational constraints. High-complexity encryption algorithms can introduce excessive computational overhead, reducing transmission frequency and accelerating battery depletion—both of which could hinder the system’s ability to respond promptly to life-critical scenarios. Another challenge is key distribution. While private key encryption enhances security, it may prevent medical professionals from accessing patient data in emergencies [5].
To address these challenges, several studies have proposed encryption methods specifically tailored for AID systems. Table 8 summarizes these approaches, detailing their methodologies, key sizes, security strengths, and limitations.
@@ 5.1.2 Encryption Techniques
Table 8: Proposed cryptography-based approaches in AID systems.
Li et al. [12] first demonstrated the vulnerability of wireless communication between glucose monitors and insulin delivery systems due to the lack of standard cryptography, they proposed using rolling code-based encryption—similar to that used in garage door systems—to prevent eavesdropping and replay attacks. However, the security of rolling code methods does not offer strong authentication and is highly dependent on encryption and decryption algorithms. Some methods, such as Keeloq, have been successfully compromised [112]. To enhance security, Marin et al. [76] employed Advanced Encryption Standard (AES), a more robust symmetric encryption algorithm. Additionally, they reduced message size and transmission frequency to optimize computational efficiency. Their experiments also revealed that authentication for messages consumed more energy than encryption due to the need to append a message authentication code (MAC), which inevitably increased message length. A more lightweight solution proposed by ForkAE [100] combined encryption and authentication processes to reduce computational overhead.
Both AES and ForkAE rely on symmetric encryption, where a public key is used for both encryption and decryption. However, this approach carries the risk of key interception. To mitigate this problem, asymmetric encryption was introduced in [70], using a private key in every packet. While this approach enhances security, it requires larger key sizes, resulting in higher computational costs and vulnerability to factorization attacks. To further protect patient privacy during computation, Weng et al. [77] proposed using homomorphic encryption within a PID-based controller. This approach enables secure data processing without key sharing, as it allows computations to be performed on encrypted data. However, the CKKS homomorphic encryption method hired in their study does not support division and comparison operations, potentially leading to imprecise computational results. Moreover, its high computational demands significantly limit its application to more complex controller algorithms.
5.1.3 Secure Communication Protocols

### v2

@@ IV Attack Vectors in AID systems
Ransomware Attacks [122] The malware encrypts the system’s data or locks users out, demanding a ransom to restore access
Jamming Attacks [112] Disrupting the wireless communication channels that the APS uses to operate
Firmware/ Software Corruption [122] Corrupting the firmware or software that the APS relies on to function
Routing protocol information attacks [19] Attacks on routing protocols to disrupt or manipulate the flow of information
TABLE V: Defence strategies proposed for AID systems.
Defence type Description Methods Protected communication Authentication Verify the identity of users or devices Fingerprint recognition[70][145]; Voiceprint analysis[106]; Physical Unclonable Functions[138][55]; Blockchain[138][110] Cryptography Transform information into an unreadable format for unauthorized users Rolling-code encryption[73]; AES-based encryption[27][81]; Asymmetric encryption [70]; Homomorphic encryption [140] Secure communication protocols Define rules for secure data transmission RF-Harvested energy-based protocol[122]; Emergencies-resilient protocol[64]; Improved the Dexcom G4 protocol[112]; Mutual authentication protocol[69] Proximity communication channels Require close physical proximity between the communicating devices Body-coupled communication[73]; Visible light communication[144] Intrusion detection systems Signature-based Compare the real-time behaviour of the system against known security attacks Define computational, data integrity, and communication attacks[62]; Define abnormal signal’s physical and content features[143]; Define replay and DoS attacks[137]; Infusion site failures [49]; Pressure-induced sensor attenuations [16] Specification-based Define specifications of AID systems’ normal operations and deem any violation as an intrusion Formal specification [25] with personalized thresholds[146][134]; statistical analysis[10]; declarative property [62]; outlier detection[13][11]; specification mining[9]; external and internal state[107][102]; insulin predictive model[134] Anomaly-based Identify data patterns that have different data characteristics from normal instances Model-based methods:
@@
Control strategies assessment Compare the dosage with the identified patient’s insulin infusion pattern Rule-based [62]; Data-driven[44][72]; Hybrid[146] Redundant hardware or software components Use of multiple, often independent, components that perform the same function Hardware: ECG monitor[102]; Universal Software Radio Peripheral[143]; Separate computation core[122]
@@
Software: Replay attack detection filters[97]; Auxiliary safety layer for controllers[113]
IV-A Threats to Confidentiality
Ensuring the confidentiality of diabetes devices aims to guarantee that sensitive information is accessible only to authorised users or systems. Threats to this security aspect can originate from unauthorised individuals or even from patients themselves.
IV-A1 Eavesdropping
This refers to the unauthorised interception of wireless communication, which can expose personal patient information related to therapy and physical condition, device information, and the device’s personal identification number (PIN). Laboratory experiments have shown that data transmitted by a commercial insulin delivery system without cryptography can be intercepted using off-the-shelf Universal Software Radio Peripheral (USRP) and free software [73][81]. By capturing its identification number, attackers could then gain full control of the pump. At a 2012 security conference, an engineer showed that a fatal dose of insulin could be delivered from 300 feet away [104]. These attacks exploit wireless communication vulnerabilities, such as wide transmission range and inadequate encryption. To mitigate these risks, countermeasures proposed in the literature mainly focus on (1) enhancing data protection through cryptography and secure communication protocols, and (2) limiting transmission range by employing proximity communication channels.
IV-A2 Do-it-yourself (DIY) hacking
This refers to unauthorised access or modification of commercial diabetes devices by patients themselves, driven by the desire for better data visualisation and more advanced control with shorter official censor and approval time [66]. For instance, Nightscout is an open-source project that helps patients hack their Dexcom CGMs for remote glucose level monitoring on mobile devices. Projects such as #DIYPS enable patients to control insulin delivery. While these initiatives empower patients, they introduce risks due to vulnerabilities in unauthorised software and users’ potential lack of expertise. Compared to FDA-regulated medical devices, these systems lack rigorous testing and responsible parties to support them in case of problems. To balance the need for patients to have greater control with secure AID systems against malicious exploitation, it is crucial to implement robust authentication strategies.
IV-B Threats to Integrity
The integrity of AID systems ensures that their data, algorithms, and operations remain accurate and consistent. Integrity can be compromised through two key aspects: data manipulation and algorithm manipulation.
IV-B1 Manipulation of Data
Attackers can manipulate wirelessly transmitted data, such as CGM readings or controller commands, to mislead the control algorithm or insulin pumps into issuing incorrect insulin delivery. Data manipulation can occur through the following attack vectors:
  • Replay attacks: These attacks involve intercepting and replaying legitimate data packets to manipulate insulin delivery [97][85]. This type of attack is possible when systems lack freshness checks or authentication, and attackers can observe and record communications. It does not require device access or message decryption.
Once the attacker gains access to the device (e.g., by accessing the PIN) and decrypts the message, they can craft forged data packets to be accepted by the device [73], enabling bias injection and false data injection attacks.
  • Bias Injection Attacks: In this attack, the attacker skews sensor readings by injecting a bias [126]. By making subtle, consistent changes, the attacker can cause the system to under- or over-deliver insulin without needing extensive knowledge of the patient model or system operations. They can be difficult to detect due to their subtlety and correlation with the original data [123].
  • False Data Injection Attacks (FDIA): FDIA refers to an adversary deliberately manipulating measurement or communication data to affect system state estimation or control decisions. They can be broadly categorised into random and targeted types, depending on how the data is forged. Random FDIA introduces arbitrary perturbations into input data. Experiments in [60] showed that such attacks can shift the outputs of deep learning (DL) based glucose prediction models by 10-20 mg/dL under different scenarios. These attacks are generally easier to detect using sequential change detection techniques.
    In contrast, targeted FDIA carefully crafts input values to evade alarms and cause deliberate harm. This belongs to process-aware attacks that require knowledge of the control algorithm, which is often detailed in documentation. These attacks can be divided into model-specific and patient-specific variants. Model-specific FDIA exploit vulnerabilities in the target model; for example, Elnawawy et al. [35] used URET to craft semantically and functionally plausible adversarial data points for interference AI models, while Kashyap et al. [60] formulated this attack as an optimisation problem, showing that minimal, hard-to-detect perturbations against DL-based models can be computed within 60 milliseconds. Patient-specific FDIA endangers selected individuals, such as personalised insulin dose manipulation attacks craft attacks that harmful for target patients and within safety constraints [72], register false states (e.g., meal intake) to covertly mislead the control logic [121], or remote control insulin delivery [73][27][81].
Defence strategies for these data manipulation attacks typically include developing intrusion detection systems to detect abnormal patterns in sensor data and assessing controller activities. Another defensive measure involves using additional sensors or software monitors to cross-check data sources for inconsistencies.
IV-B2 Manipulation of Algorithms
Algorithm manipulation involves tampering with the computational logic or parameter settings of AID systems, targeting either control algorithms or physiologic models used for patient state prediction [62]. These include both traditional rule-based algorithms and machine learning(ML)-based models.
  • Computational attacks: These attacks may compromise instructions or internal data in algorithm execution [62]. For example, Aliabadi et al. [8] identified that the synchronisation between CGM and controller could be manipulated by altering the timing function (get-data-timer()). This modification allows attackers to delay or drop critical messages. Other threats involve tampering with essential device settings, such as disabling pump alarms, modifying the maximum dosage limits [27] or pump I/O operations [134].
  • ML model attacks: ML models are increasingly used in AID systems to enhance performance [53]. For instance, the FDA-approved One Drop Blood Glucose Monitoring System employs a multi-layer perceptron (MLP) model for glucose prediction. Deep reinforcement learning (RL) approaches have been explored for use in AID systems for BG control [39]. However, due to their black-box and data-dependent nature, understanding and monitoring the reasoning behind these models is challenging. Attackers may influence the model’s behaviour during training or inference phases or reverse-engineer the model to extract confidential information [35]. Moreover, ML-based algorithms can be severely interfered with by a very slight input data perturbation that is hard to detect—an attack variant described as ripple FDIA [60]. In [60], minimal interference was calculated to misguide the ML model into predicting a patient’s glucose level as hyperglycaemic, leading the controller to administer an excessive dose of insulin. Experiments in [24] show that a small perturbation around ±1mg/dL on BG reading can lead to severe patient harm over time in an RL controller model.
To defend against algorithm integrity threats, formal verification techniques are often employed to assess the operations of rule-based algorithms. For ML-enabled systems, where the black-box nature poses challenges for formal verification, intrusion detection systems and modelling insulin injection patterns can be used to detect abnormalities and inconsistencies.
IV-C Threats to Availability
  • Denial-of-Service (DoS) and Distributed DoS (DDoS): In these attacks, an adversary floods the system with excessive traffic, overwhelming AID systems and preventing legitimate requests from being processed promptly [134]. This could lead to delays or complete failure in delivering insulin [137].
  • Ransomware Attacks: Ransomware can encrypt data and lock critical components of AID systems, demanding ransom payments to restore access [122]. This can prevent patients or healthcare providers from viewing essential data, such as glucose levels or delivery history, rendering the system inoperable until resolved.
  • Jamming Attacks: By interfering with the wireless communication frequency, attackers can interrupt data flow, causing the system to fail to deliver insulin as required [112].
  • Firmware/Software Corruption: Malicious firmware or software corruption can cause AID systems devices to malfunction or stop functioning entirely [122]. For example, saturation-based sensor spoofing attacks make a sensor ignore legitimate inputs by using an additional infrared source to saturate the sensors [103]. Additionally, attacks may exploit software vulnerabilities or delivering harmful updates can prevent systems from operating as intended, leading to a cessation of insulin administration.
  • Routing Protocol Information Attacks: Manipulating data transmission pathways in networked AID systems, attackers can disrupt or delay communication between key system components, compromising timely insulin delivery and glucose monitoring [19].
Mitigation strategies for these threats include network traffic monitoring, secure communication protocols, authentication mechanisms, formal verification, and redundant hardware or software components.
V Defence Strategies for AID systems
Addressing the wide range of potential cyberattacks on AID systems requires comprehensive defence strategies that span multiple system layers. Defence mechanisms in AID systems can be grouped into three primary areas: protected communication, intrusion detection systems, and control strategies assessment. Besides these internal defences, external strategies—such as the use of redundant hardware or software components—are employed to detect abnormal activities independent of AID devices. Table V provides an overview of the defence types and corresponding methods proposed in past research.
V-A Protected Communication
@@ V-A1 Authentication Methods
Biometric authentication methods like fingerprint recognition [70][145] and voiceprint analysis [106] are user-friendly and offer strong uniqueness. However, they require active user intervention, making them more appropriate for infrequent control operations (e.g., bolus injections) rather than continuous biosignal transmissions that occur periodically (e.g., every 5 minutes). Moreover, they raise concerns regarding the secure storage of sensitive biometric data, which could be spoofed by advanced attacks. To enhance system autonomy and reduce the risk of key leakage, hardware-based security methods such as Physically Unclonable Functions (PUFs) exploit inherent and unique physical variations in manufacturing processes, offering low computational overhead and an unclonable security mechanism without the need for key storage [138][55]. Additionally, blockchain technology has been explored to protect patient data by decentralising data storage in central cloud servers [138]. Rahmadika et al. [110] further incorporated blockchain with federated learning for sharing local parameters for the misbehaviour detection model in AID systems without compromising patient privacy.
V-A2 Encryption Techniques
Following successful authentication, establishing encrypted communication between AID systems components is vital for maintaining data confidentiality and integrity. A key challenge is the energy and computational constraints. High-complexity encryption algorithms can introduce excessive computational overhead, reducing transmission frequency and accelerating battery depletion—which could hinder the system’s ability to respond promptly to critical scenarios. Another challenge is key distribution. While private key encryption enhances security, it may prevent medical professionals from accessing patient data in emergencies [143].
To address these challenges, several studies have proposed encryption methods specifically tailored for AID systems. Table VI summarises these approaches, detailing their methodologies, key sizes, security strengths, and limitations.
TABLE VI: Proposed cryptography-based approaches in AID systems.
@@ V-A2 Encryption Techniques
Li et al. [73] first demonstrated the vulnerability of wireless communication between glucose monitors and insulin delivery systems due to the lack of standard cryptography; they proposed using rolling code-based encryption—similar to that used in garage door systems—to prevent eavesdropping and replay attacks. However, the security of rolling code methods does not offer strong authentication and is highly dependent on encryption and decryption algorithms. Some methods, such as Keeloq, have been successfully compromised [96]. To enhance security, Marin et al. [81] employed Advanced Encryption Standard (AES), a more robust symmetric encryption algorithm. Additionally, they reduced message size and transmission frequency to optimize computational efficiency. Their experiments also revealed that authentication for messages consumed more energy than encryption due to the need to append a message authentication code (MAC), which inevitably increased message length. A more lightweight solution proposed by ForkAE [27] combined encryption and authentication processes to reduce computational overhead.
Both AES and ForkAE rely on symmetric encryption, where a public key is used for both encryption and decryption. However, this approach carries the risk of key interception. To mitigate this problem, asymmetric encryption was introduced in [70], using a private key in every packet. While this approach enhances security, it requires larger key sizes, resulting in higher computational costs and vulnerable to factorisation attacks. To further protect patient privacy during computation, Weng et al. [140] proposed using homomorphic encryption within a PID-based controller. This approach enables secure data processing without key sharing, as it allows computations to be performed on encrypted data. However, the CKKS homomorphic encryption method used in their study does not support division and comparison operations, potentially leading to imprecise computational results. Moreover, its high computational demands limit its application to more complex controller algorithms.
V-A3 Secure Communication Protocols